# A Novel Spike Transformer Network for Depth Estimation from Event Cameras via Cross-modality Knowledge Distillation

### v2

@@ 2.3 Knowledge distillation for SNN
[…]
However, existing knowledge distillation methods for SNNs all require training a teacher model first. Currently, large foundation models become the new deep learning hotspot [47].A large foundation model is trained on a vast quantity of data at scale (often by self-supervised learning or semi-supervised learning) so that the learned features can be used directly for various downstream tasks or knowledge distillation. For example, Dense Prediction Transformers (DPT) [32] are a type of ViT that is designed for depth prediction tasks, which was trained on 1.4 million images for monocular depth estimation. DINOv2 [30] used ViT-Giant, a larger version of ViT with 1 billion parameters. It is more powerful than previous ViT models and outperforms previous self-supervised learning methods in a variety of computer vision tasks, especially for depth estimation. In this work, for the first time, we will explore to transfer knowledge from a large foundation model (DINOv2) [30]to SNNs for depth estimation.
@@ 3 The Proposed Method
We introduce a novel pure spike-based transformer computational network for depth estimation via cross-modality knowledge distillation. The network architecture employs an encoder-decoder structure, with the transformer serving as the core computational unit of the encoder. A fusion decoder is specifically crafted to amalgamate features from different stages of the encoder, thus producing a comprehensive depth map. The flowchart of the method is illustrated in Figure 1, encompassing three primary components: 1) Spike-driven transformer, 2) Knowledge distillation, and 3) Fusion depth estimation head.
The underpinnings of our proposed methodology are outlined as follows:
  1. 1.
    Transformers, renowned for their capability to capture long-range dependencies through self-attention and matrix multiplications, typically require computationally intensive operations that are not well-suited for the spike-driven paradigm. In this context, we introduce a pure spike-driven transformer network that leverages spike-driven residual learning and spike self-attention. This design eschews conventional floating-point and integer-float multiplications, adhering strictly to spiking principles and thus achieving substantial energy savings while maintaining robust performance.
  2. 2.
    SNNs are energy-efficient models that use binary spikes instead of continuous values, but they also need deep architectures to achieve high accuracy, which can be computationally expensive and power-hungry. Knowledge distillation allows leveraging the knowledge of a pre-trained ANN (teacher) to guide the training of a smaller SNN (student) model, enabling it to achieve similar performance with fewer resources. In this work, a cross-modality knowledge distillation method is proposed between RGB and event data. The proposed method is trained through matching RGB and event data, while the inference is performed on event data only. The large pretrained vision foundation model DINOv2 is used as the teacher model for knowledge distillation, which does not participate in training and only extracts features from the RGB data.
  3. 3.
    Differing from conventional fully convolutional networks that employ downsampling, transformer networks maintain consistent dimensional representations, facilitating global receptive fields essential for tasks such as segmentation or depth prediction. Hence, we propose a fusion depth estimation head capable of harnessing features from each transformer stage to optimize depth estimation effectiveness.
3.1 Pure spike-driven transformer network for depth estimation
In this work, we propose a pure spike transformer-computational network for depth estimation via cross-modality knowledge distillation. The design of the proposed method follows a pattern that logically divides the network into an encoder and a decoder. The spike-transformer block is selected as the encoder to extract spatial-temporal features from spike data. A fusion depth estimation head is designed as the decoder to generate pixel-scale depth results from the feature representation.
3.1.1 Spike-transformer
@@
The proposed spike transformer aligns with the foundational structure of the original Vision Transformer (ViT), encompassing a Spiking Patch Embedding and Spiking Transformer Block. Given an event sequence. $I\in\mathbb{R}^{T\times C\times H\times W}$, the spike patch embedding is used to convert the input into a sequence of tokens that can be processed by the transformer architecture, where the event input is projected as spike-form patches $X\in\mathbb{R}^{T\times N\times D},N=\frac{H}{8}\times\frac{W}{8}$. Then, the spiking patches $X$ are passed to the multi spiking transformer blocks (L). Considering that we have used knowledge distillation from the large model, this method uses only a minimum number of blocks as $\mathrm{L}=4$. Inspired by [27, 28], in order to avoid non-spike computations in traditional deep learning architectures, a Spiking Self Attention (SSA) and a Spiking MLP block are used in spiking transformer blocks.
Spiking patch embedding
@@
In the original ViT [20], the patch embedding is used to represent an image as a sequence of tokens. This is done by dividing the image into a grid of patches and flattening each patch into a vector. In this work, we implement this operation through a convolution batch norm (ConvBN), Max pooling (MP) and multistep LIF (MLIF) combination. The structure is shown in Figure 2. This process can be formulated as:
@@
Spiking Transformer Block
@@
Guided by the findings in [28], we position a multistep LIF before the ConvBN within the residual mechanism to omit floating-point multiplication and mixed-precision calculations during the ConvBN operation. This adjustment also enables ConvBN to replace conventional linear layers and batch normalization seamlessly. The SSA operation can be mathematically described as:
@@
The variables $Q,K,V\in\mathbb{R}^{T\times N\times D}$ represent pure spike data (only containing 0 and 1). The Scaling factor, s, is used to adjust the largest value of the matrix multiplication result. It does not affect the property of SSA. The Spiking MLP block consists of a residual connection and a combination of MLIF and ConvBN.
3.1.2 Fusion Depth estimation Head
The aim of depth estimation head is to predict the depth for each pixel which is similar to the segmentation task. The most dominant segmentation framework currently is the U-Net based multiscale structure. The decoder of U-Net uses skips connections from layers of different scale in the encoder to help preserve the spatial information of input images. This helps to ensure that the decoder has access to both the low-level and high-level features from the encoder, in which the low-level features are important for capturing details and the high level features are important for understanding the overall structure of the image. Unlike traditional CNN structures, the size of hidden features in the transformer structure is invariant and it is not straightforward to use the U-Net structure. There are a number of models that use transformer as the backbone for segmentation and redesign the structure of the transformer, such as Multiscale vision transformers [48], Pyramid vision transformer [49], Multi-Path Vision Transformer [50], etc. In this work, a knowledge distillation is selected to distil the knowledge from the trained large transformer into the proposed SNN. Modifying the structure will significantly reduce the effectiveness of the distillation. Therefore, a fusion depth estimation head is proposed to predict the depth.
@@ 3.1.2 Fusion Depth estimation Head
The structure of the fusion head for depth estimation is shown in Figure 4. The first step of the fusion head is to assemble the internal features in transformer blocks into image-like feature representations. The feature representations are then fused into the final dense prediction with skip connections. A generic up sample structure is used to restore the feature representations to original data size. Give an input feature as $F_{i}\in\mathbb{R}^{T\times H/8\times W/8}$), i$=$1,2,3,4. The depth estimation head can be formulated as follows:
$\begin{gathered}Y=\operatorname{ConvBn}\left(Up\left(\operatorname{ConvBn}\left(\operatorname{ConvBn}\left(\operatorname{ConvBn}\left(\left(\operatorname{ConvBn}\left(Up\left(F_{1}\right)\right)+Up\left(F_{2}\right)\right)\right.\right.\right.\right.\right.\\ \left.\left.\left.\left.+Up\left(F_{3}\right)\right)+Up\left(F_{i+4}\right)\right)\right)\right)\\ Y=\operatorname{Sigmod}(Y)\end{gathered}$ (4)
3.2 Knowledge distillation from DINOv2
@@
In this work, a method of cross-modality knowledge distillation is outlined, wherein a large-scale vision foundation model, specifically DINOv2, is utilized to guide the training of our SNN. The DINOv2 model, trained on a diverse dataset comprising 142 million images [30], is advantageous for several reasons:
  1. 1.
    The architecture of DINOv2, based on the Vision Transformer (ViT), resembles that of our model, facilitating effective knowledge distillation owing to their structural and feature-size compatibility.
  2. 2.
    The features extracted from the DINOv2 pre-trained model have been used to achieve the state-of-the-art performance on the NYU and SUN RGB-D depth estimation benchmark datasets. Figure 5 shows the visualization of self-attention on DIVO_V2 features and the depth estimation result with a linear probe on frozen DINOv2 features on our dataset. The DIVNO v2’s feature representations (the middle column in Figure 5) and the depth estimation result close to the real depth of the image, as visualized in the figure, supports the hypothesis that the DINO V2 model can provides effective guidance during the knowledge distillation training process.
Figure 1 shows the knowledge distillation process. We freeze the DINOv2 as a teacher model. The output features from DINOv2 are considered as targets in our training. To obtain the same feature size, we upsample the RGB image by a factor of 1.75. The final feature size of teacher model is $x_{rgb}^{{}^{\prime}}\in\mathbb{R}^{d\times H/8\times W/8}$.
@@
Figure 6 shows the knowledge distillation algorithm. In our approach, we utilize a fusion loss function. The L1 loss measures the differences in features between the student and teacher networks. The Perceptual Loss [51], a method commonly used in machine learning to quantify discrepancies between two images, is also adopted here. The L2 loss function, a scale-invariant metric [52], is chosen to evaluate the results of depth estimation. This scale-invariant feature ensures that the loss does not penalize the model for discrepancies arising from the scale of input data, demonstrating effectiveness for monocular depth estimation. The scale-invariant loss is defined by the following equations:
$\begin{gathered}\mathcal{L}_{1}=\frac{1}{C\times H\times W}\left\|x-x^{\prime}\right\|_{2}^{2}\\ \mathcal{L}_{2}=\frac{1}{n}\sum_{\mathbf{i}}\left(D_{i}^{t}-D_{i}^{p}\right)^{2}-\frac{1}{n^{2}}\left(\sum_{\mathbf{p}}D_{i}^{t}-D_{i}^{p}\right)^{2}\end{gathered}$ (5)
Where $D_{i}^{t}-D_{i}^{p}$ is the difference between predicted and ground truth depth for pixel i,and n is total number of pixels with a dimension of $H\times W$.
@@ 4.1 Datasets
The first dataset is a synthetic dataset from [26], which is generated from DENSE dataset [Hidalgo-Carrió2020Learning], including clear depth maps and intensity frames in 30 FPS under a variety of weather and illumination conditions. To obtain spike streams with high temporal resolution, the video is interpolated to generate intermediate RGB frames between adjacent 30-FPS frames. With absolute intensity information among RGB frames, each sensor pixel can continuously accumulate the light intensity with the spike generation mechanism, producing spike streams with a high temporal resolution (128$\times$30 FPS) that is 128 times of the video frame rate. The ‘spike’ version of DENSE dataset (namely DENSE spike) contains eight sequences, five for training, and three for evaluation. Each sequence consists of 999 samples, and each sample is a tuple of one RGB image, one depth map, and one spike stream. Each spike stream is simulated between two consecutive images, generating a binary sequence of 128 spike frames (with a size of 346 $\times$ 260 each) to depict the continuous process of dynamic scenes.
@@ 4.2 Experiment design
4.2.1 Experiment 1. Qualitative and Quantitative Comparaisons
In this section, we evaluate the depth estimation performance and energy consumption of our SNN on the synthetic (DENSE) and real datasets (DSEC) and compare it with three competing dense prediction networks, namely U-Net [36], E2Depth [Hidalgo-Carrió2020Learning] and Spike-T [26]. U-Net employs 2D convolutional layers as its encoder and focuses on spatial feature extraction, while E2Depth applies ConvLSTM layers that combine CNN and LSTM to capture the spatial and temporal features. The Spike-T employs transformer-based blocks to learn the spatio-temporal features simultaneously. These models therefore constitute our immediate and direct competitors.
4.2.2 Experiment 2 Ablation study: the contribution of the proposed modules

### v3

@@ 2.3 Knowledge distillation for SNN
Knowledge distillation is a type of model compression technique that transfers knowledge from large teacher models to smaller student models. It has gained attention due to its ability to train deep neural networks using limited resources (Gou et al., 2021), and has been shown effective in improving the performance of SNNs. In Kushawaha et al. (2020), the author proposed a knowledge distillation method for transferring knowledge from a large trained SNN to a small one in an image classification task. The results show that using the knowledge of pre-trained large models can significantly improve the performance of small models, thus enhancing the possibility of deploying high-performance models on resource-limited platforms. In study (He et al., 2023), a knowledge distillation method was proposed for SNNs performing image classification. The method was able to improve the accuracy of a student SNN by 2.7% to 9.8%. The authors (Qiu et al., 2022) found that the knowledge distillation paradigm can be effective in reducing the performance gap from the ANN to the SNN. The distillation between ANNs and SNNs using similar structures improved model performance. In (Liu et al., 2022a), knowledge distillation was first used in training SNNs for depth estimation. The author proposed a cross-modality domain knowledge transfer method for unsupervised spike depth estimation with open-source RGB data.
While knowledge distillation has shown promise in improving SNN performance (He et al., 2023; Qiu et al., 2022), existing approaches have several limitations: 1) Most methods focus on classification tasks and cannot be directly applied to dense prediction problems like depth estimation; 2) Cross-modality knowledge transfer between conventional images and event data remains largely unexplored; 3) Current approaches typically require training a separate teacher model, increasing computational overhead. Currently, large foundation models have become the new deep learning hotspot (Bommasani et al., 2021). A large foundation model is trained on a vast quantity of data at scale (often by self-supervised learning or semi-supervised learning) so that the learned features can be used directly for various downstream tasks or knowledge distillation. For example, Dense Prediction Transformers (DPT) (Ranftl et al., 2021) are a type of ViT designed for depth prediction tasks, trained on 1.4 million images for monocular depth estimation. DINOv2 (Oquab et al., 2023) used ViT-Giant, a larger version of ViT with 1 billion parameters. It is more powerful than previous ViT models and outperforms previous self-supervised learning methods in a variety of computer vision tasks, especially for depth estimation. In this work, for the first time, we will explore transferring knowledge from a large foundation model (DINOv2) to SNNs for depth estimation.
@@ 3 The Proposed Method
We propose a novel energy-efficient spike transformer network for depth estimation via cross-modality knowledge distillation. The flowchart of the method is illustrated in Figure 1, encompassing three primary components: 1) Spike-driven transformer, 2) Fusion depth estimation head, and 3) Knowledge distillation. The rationale for our proposed method centres on three key innovations:
1) We introduce a pure spike-driven transformer network that eliminates conventional floating-point operations. While transformers excel at capturing long-range dependencies, their matrix multiplication operations are computationally intensive. Our spike-driven design using binary spikes instead of floating-point values significantly reduces energy consumption while maintaining high performance through carefully designed spike-based attention and residual mechanisms.
2) We propose a novel fusion depth estimation head designed to integrate features from multiple transformer stages for precise and robust depth estimation. Compared to existing methods, our approach overcomes the limitations of CNN based architectures, which often lose critical spatial information due to downsampling. By leveraging transformers’ ability to retain dimensional consistency and integrating features at multiple levels, the fusion head achieves superior depth estimation accuracy. Additionally, it is fully compatible with spike-based computation models, making it both efficient and biologically plausible, providing a significant advantage for real-world applications requiring precision and robustness in challenging environments.
3) To address the limited training data available for SNNs, we leverage knowledge from DINOv2, a large vision foundation model, through a novel single-stage cross-modality distillation framework. Rather than requiring separate training phases or an additional teacher model, our approach directly transfers relevant features from RGB to event data domains, enabling efficient training while preserving the spike-based computation paradigm.
3.1 Spike-driven transformer
@@
The proposed spike transformer aligns with the foundational structure of the original Vision Transformer (ViT), encompassing a Spiking Patch Embedding and Spiking Transformer Block. Given an event sequence, $I\in\mathbb{R}^{T\times C\times H\times W}$, the spike patch embedding is used to convert the input into a sequence of tokens that can be processed by the transformer architecture, where the event input is projected as spike-form patches $X\in\mathbb{R}^{T\times N\times D},N=\frac{H}{8}\times\frac{W}{8}$. Then, the spiking patches $X$ are passed to the multi spiking transformer blocks (L). Considering that we have used knowledge distillation from the large model, this method uses only a minimum number of blocks as $\mathrm{L}=4$. Inspired by (Zhou et al., 2022, 2023), in order to avoid non-spike computations in traditional deep learning architectures, a Spiking Self Attention (SSA) and a Spiking MLP block are used in spiking transformer blocks.
3.1.1 Spiking patch embedding
@@
In the original ViT (Dosovitskiy et al., 2020), the patch embedding is used to represent an image as a sequence of tokens. This is done by dividing the image into a grid of patches and flattening each patch into a vector. In this work, we implement this operation through a convolution batch norm (ConvBN), Max pooling (MP) and multistep LIF (MLIF) combination. The structure is shown in Figure 2. This process can be formulated as:
@@
3.1.2 Spiking Transformer Block
@@
Guided by the findings in (Zhou et al., 2023), we position a multistep LIF before the ConvBN within the residual mechanism to omit floating-point multiplication and mixed-precision calculations during the ConvBN operation. This adjustment also enables ConvBN to replace conventional linear layers and batch normalization seamlessly. The SSA operation can be mathematically described as:
@@
The variables $Q,K,V\in\mathbb{R}^{T\times N\times D}$ represent pure spike data (only containing 0 and 1).The Q, K, and V matrices are first obtained through learnable transformation matrices. These matrices are then converted into spiking sequences using distinct spiking neuron layers. The SSA leverages the inherently non-negative properties of the spike-form Q and K to produce a non-negative attention map. This makes softmax redundant as SSA can directly aggregate relevant features while ignoring irrelevant ones. The Scaling factor, s is used to adjust the largest value of the matrix multiplication result. It does not affect the property of SSA. The Spiking MLP block consists of a residual connection and a combination of MLIF and ConvBN.
3.2 Fusion Depth estimation Head
The task of depth estimation requires generating pixel-wise depth predictions from encoded features. Two common approaches exist: the Fully Convolutional Network (FCN) head and fusion-based architectures. The FCN head applies convolution operations directly to the final encoder features to predict depths. While computationally efficient, this approach relies solely on high-level features and often struggles to preserve fine spatial details.
Unlike CNN architectures that naturally produce multiscale feature hierarchies through downsampling, transformer networks maintain constant feature dimensions across layers. This poses a unique challenge for depth estimation, which benefits from both coarse semantic information and fine spatial details. Previous transformer-based segmentation models have addressed this through architectural modifications like the Multi-Scale Vision Transformer (Fan et al., 2021) or Pyramid Vision Transformer (Wang et al., 2021). However, such modifications could reduce the effectiveness of knowledge distillation from pretrained vision models because they are not modified with the same structure.
Therefore, we propose a fusion depth estimation head that leverages features from multiple transformer stages while preserving the original transformer architecture. The fusion head first projects the internal features from each transformer block into image-like representations. These multi-level features are then progressively combined through skip connections and upsampling operations to generate the final dense prediction.
The structure of the fusion head for depth estimation is shown in Figure 4.
@@ 3.2 Fusion Depth estimation Head
The first step of the fusion head is to assemble the internal features in transformer blocks into image-like feature representations. The feature representations are then fused into the final dense prediction with skip connections. A generic up sample structure is used to restore the feature representations to original data size. Give an input feature as $F_{i}\in\mathbb{R}^{T\times H/8\times W/8}$), i$=$1,2,3,4. The depth estimation head can be formulated as follows:
$\begin{gathered}\mathrm{Y}_{2}=\left(\operatorname{ConvBn}\left(\operatorname{Up}\left(F_{1}\right)\right)+\operatorname{Up}\left(F_{2}\right)\right)\\ \mathrm{Y}_{3}=\left(\operatorname{ConvBn}\left(\operatorname{Up}\left(\mathrm{Y}_{2}\right)\right)+\operatorname{Up}\left(F_{3}\right)\right)\\ \mathrm{Y}_{4}=\left(\operatorname{ConvBn}\left(\operatorname{Up}\left(\mathrm{Y}_{3}\right)\right)+\operatorname{Up}\left(F_{4}\right)\right)\\ Y=\operatorname{Sigmod}\left(\mathrm{Y}_{4}\right)\end{gathered}$ (4)
This design enables the network to combine high-level semantic information from deeper layers with fine-grained spatial details from earlier layers, leading to more accurate depth predictions while maintaining compatibility with knowledge distillation from vision foundation models.
3.3 Knowledge distillation
@@
Knowledge distillation is particularly challenging for SNNs due to two main factors: (1) the binary nature of spike data differs fundamentally from the continuous values used in traditional ANNs, and (2) the limited availability of labeled event camera data makes training challenging. To address these challenges, we propose a single-stage cross-modality knowledge distillation framework that leverages DINOv2 (Oquab et al., 2023), a large-scale vision foundation model, to guide our SNN training.
Our choice of DINOv2 as the teacher model is motivated by several key advantages:
1) Architectural Compatibility: DINOv2’s Vision Transformer (ViT) architecture closely aligns with our model’s structure, facilitating effective knowledge transfer due to their similar feature representations and computational patterns.
2) Rich Feature Representations: Pre-trained on 142 million diverse images, DINOv2 has demonstrated state-of-the-art performance in depth estimation tasks on benchmark datasets like NYU and SUN RGB-D. As shown in Figure 5, DINOv2’s self-attention patterns and depth estimation capabilities on our dataset suggest it can provide valuable guidance during the knowledge distillation process.
3) Zero-shot Generalization: DINOv2’s strong zero-shot learning capabilities enable effective knowledge transfer even when dealing with limited event camera data.
The knowledge distillation process can be shown in Figure 1. We freeze the DINOv2 (Lightblue) as a teacher model. The output features from DINOv2 are considered as targets in our training. To ensure compatible feature dimensions, we upsample the RGB input images by a factor of 1.75, resulting in teacher model features of size $x_{rgb}^{{}^{\prime}}\in\mathbb{R}^{d\times H/8\times W/8}$.
@@
Figure 6 shows the knowledge distillation algorithm. Our distillation framework employs a fusion loss function that combines two complementary components:
Feature Perceptual Loss (Lp): Measures the distance between student and teacher feature representations, ensuring the SNN learns similar feature patterns. The Perceptual Loss (Johnson et al., 2016) is used here to help capture high-level semantic differences between teacher and student representations, going beyond pixel-level comparisons.
The L2 loss function, a scale-invariant metric (Eigen and Fergus, 2015), specifically designed for monocular depth estimation to address the inherent scale ambiguity problem.This scale-invariant loss is particularly important as it focuses on relative depth relationships rather than absolute values, aligning with the fundamental nature of monocular depth estimation where absolute scale cannot be determined from a single view. The combination of these loss terms enables our student network to learn both detailed features and consistent global depth relationships from the teacher model while maintaining computational efficiency through spike-based operations.
The fusion loss function is defined by the following equations:
$\begin{gathered}\mathcal{L}_{Pi}=\frac{1}{C\times H\times W}\left\|x_{i}-x_{i}^{\prime}\right\|_{2}^{2}\\ \mathcal{L}_{2}=\frac{1}{n}\sum_{\mathbf{i}}\left(D_{i}^{t}-D_{i}^{p}\right)^{2}-\frac{1}{n^{2}}\left(\sum_{\mathbf{p}}D_{i}^{t}-D_{i}^{p}\right)^{2}\end{gathered}$ (5)
Where $D_{i}^{t}-D_{i}^{p}$ is the difference between predicted and ground truth depth for pixel i,and n is total number of pixels with a dimension of $H\times W$. It makes the loss invariant to uniform scaling of the depth predictions, allowing the network to learn consistent relative depth relationships even when absolute scale cannot be determined. This aligns with human depth perception, which relies heavily on relative rather than absolute depths.
@@ 4.1 Datasets
DENSE Datasets: The first dataset is a synthetic dataset from (Zhang et al., 2022), which is generated from DENSE dataset (Hidalgo-Carrió et al., 2020), including clear depth maps and intensity frames in 30 FPS under a variety of weather and illumination conditions. To obtain spike streams with high temporal resolution, the video is interpolated to generate intermediate RGB frames between adjacent 30-FPS frames. With absolute intensity information among RGB frames, each sensor pixel can continuously accumulate the light intensity with the spike generation mechanism, producing spike streams with a high temporal resolution (128$\times$30 FPS) that is 128 times of the video frame rate. The ‘spike’ version of DENSE dataset (namely DENSE spike) contains eight sequences, five for training, and three for evaluation. Each sequence consists of 999 samples, and each sample is a tuple of one RGB image, one depth map, and one spike stream. Each spike stream is simulated between two consecutive images, generating a binary sequence of 128 spike frames (with a size of 346 $\times$ 260 each) to depict the continuous process of dynamic scenes.
@@ 4.2 Experiment design
4.2.1 Model performance
In this section, we evaluate the depth estimation performance and energy consumption of our SNN on the synthetic (DENSE) and real datasets (DSEC) and compare it with three competing dense prediction networks, namely U-Net (Ronneberger et al., 2015), E2Depth (Hidalgo-Carrió et al., 2020) and Spike-T (Zhang et al., 2022). U-Net employs 2D convolutional layers as its encoder and focuses on spatial feature extraction, while E2Depth applies ConvLSTM layers that combine CNN and LSTM to capture the spatial and temporal features. The Spike-T employs transformer-based blocks to learn the spatio-temporal features simultaneously. These models therefore constitute our immediate and direct competitors.
4.2.2 Ablation study